Two children with mass 35kg slide down an icy hill on a 21kg sled. assuming not friction, what is the acceleration of the sled if the slope of the hill is 21 degrees?

Ceres is the largest asteroid in our solar system with a mass of 7 x 10^20 kg and a radius of 500km. what would a 75kg man weigh if standing on Ceres?
Answer #1

1)

angle of incline , theta = 21 degree

acceleration of the sled = g * sin(theta)

acceleration of the sled = 9.8 * sin(21 degree)

acceleration of the sled = 3.51 m/s^2

the acceleration of the sled is 3.51 m/s^2

2) let the weight of the man is W

W = G * M * m/R^2

W = 6.673 *10^-11 * 7 *10^20 * 75/(500 *10^3)^2

W = 14.01 N

the weight of the man is 14.01 N